### v1

# Made-to-Measure Outperforms Schwarzschild’s Method

Abstract
Syer and Tremaine’s made-to-measure method and Schwarzschild’s orbit superposition method are well-known within the field of stellar dynamical modeling. This research is concerned with assessing and comparing the operational capabilities of the two methods and, in particular, the impact on observable reproduction, orbit classifications and computer elapsed times when using low orbit numbers (8000 orbits) with different observational data sets and initial conditions. Both methods are able to reproduce observed data with mean $\chi^{2}\approx 1$ or less. However, the made-to-measure process does so three to five times faster than the orbit superposition method, and this starts to make the made-to-measure process attractive for analysing galaxy surveys. For a given set of initial conditions, both methods produce similar orbit classifications but the orbits behind the classifications are not the same. Orbits which are common do not have the same weights. Different initial conditions result in different classifications.
keywords:
galaxies: kinematics and dynamics – galaxies: structure – methods: numerical
1 Introduction
Syer and Tremaine’s made-to-measure method (Syer1996) (M2M) and Schwarzschild’s orbit superposition method (Schwarz1979) (SCHW) are well-known within the field of stellar dynamical modeling. Of the two, SCHW has more citations (885) at the time of preparing this article (mid-2025) than M2M (168). M2M and SCHW have been compared previously in Long2012 where M2M and SCHW were shown to be able to produce similar estimates of galaxy mass-to-light ratios. It is now appropriate that a further comparison is made.
Three factors are relevant.
  1. 1.
    Data sizes have increased both in terms of the numbers of galaxies being surveyed and the numbers of spaxels being used to capture galaxy data, and will continue to increase. Internally, because of its orbit weighting mechanism, SCHW is not well structured to deal with the increases expected. M2M is much better positioned.
  2. 2.
    Galaxy surveys would benefit from being analyzed by methods such as M2M and SCHW with their abilities to handle data from multiple surveys, data of different types (spaxel and discrete), chemical data, population synthesis data and orbit classification constraints. However both M2M and SCHW appear to be too slow to handle survey scale activities.
  3. 3.
    An improved understanding of what methods such as M2M and SCHW are capable of doing, or not doing. For example, M2M and SCHW can only weight the orbits that have been provided to reproduce the observed data. At best, the models produced may only be illustrative of what the real galaxy might be like. At worst, the models are non-physical.
This article will not address the third factor: this may be better left to machine learning. The first two factors will be considered further. At the time of the first comparison, the accepted assessment was that M2M, with its ability to use significantly more orbits (numbers of $10^{5}$ to $10^{6}$ were not uncommon), must of course be producing the better models. SCHW by comparison was claimed to model real galaxies with only a few hundred orbits. The question being posed in this article is how does M2M compare with SCHW when run using fewer orbits, for example $\approx 10^{4}$. The objective of this research therefore is to assess and compare the operational capabilities of M2M and SCHW, in particular the impact on observable reproduction, orbit classifications and computer elapsed times, when using low orbit numbers.
The article’s structure is as follows. Section 2 summarizes at a top level the approach taken in the investigation. The observational data are described in Section 3, with methods and techniques in Section 4. Results and subsequent discussion are in Sections 5 and 6, with conclusions in Section 7. The main abbreviations used are defined in Table 1.
Table 1: Abbreviations
Abbreviation Meaning
3I three integral initial conditions
ATLAS3D relating to the ATLAS3D survey
GPU graphics processing unit
hmdtu half mass dynamical time unit (galaxy specific time unit)
IFU integral field unit
M2M relating to the made-to-measure method
MDJV ‘match density Jeans velocities’ initial conditions
MGE multi-Gaussian expansion
MPI message passing interface
M/L mass-to-light
NNLS non-negative least squares
PC personal computer
SCHW relating to Schwarzschild’s method
2 Approach
The Introduction, Section 1, sets the context and describes what is to be investigated. This Approach section describes how the investigation is performed. In this case, the approach is straightforward: it is just a simple combinatorial problem of running a number of models using different methods, initial conditions, and data sets. The results required for comparison purposes are
  • $\bullet$
    the observable mean $\chi^{2}$ values indicating how well the model has reproduced its observational data,
  • $\bullet$
    the computer elapsed time for the modeling runs, and
  • $\bullet$
    the orbit classifications giving the weighted hot, warm, cold and counter-rotating orbit fractions supporting reproduction of the observables.
The assessment of the performance of M2M relative to SCHW will be based on these comparisons. Gravitational potentials and initial conditions are specific to the observational data, and are only calculated once since they will be reused whenever the observational data are used. Data layouts (for example, using a polar grid or Voronoi cells) are not changed except for one set of very specific tests. For other parameters, such as the duration of model runs, their values are kept the same for both the M2M and SCHW runs.
In outline, there are
  • $\bullet$
    two methods, M2M and SCHW;
  • $\bullet$
    two sets of initial conditions, theoretically based (3I) and observationally based (MDJV); and,
  • $\bullet$
    two top level data sets, real galaxy data ( ATLAS3D ) and simulated galaxy data (IllustrisTNG).
More information is provided in Sections 3 and 4.
It is not intended that the behavior and performance of all the capabilities of M2M will be compared with all the capabilities of SCHW but only the core capabilities concerned with luminosity and kinematical data reproduction. Features such as discrete data handling, or chemical modeling, or orbit classification constraints are out of scope, as is parameter estimation using the methods. Note also that only axisymmetric (oblate) models are utilized: the investigations are not affected by choice of gravitational potential. The kinematical observables are varied so that both velocity moments and Gauss-Hermite coefficients are modeled - see Section 3. The earlier comparison in Long2012 used Gauss-Hermite coefficients only.
SCHW requires the use of NNLS software to create orbit weights, We choose NNLS software that implements the Lawson and Hanson active set method (LH1974) since it appears from the literature to be the most used NNLS implementation for SCHW. The active set algorithm deliberately zeroises the weights of any orbits it chooses not to use in reproducing the observed data, and this is taken into account in the analysis software used.
Model selection frameworks such as Lipka2021 and Thomas2022 are not used. Some use is made of an experimental Gaussian processes approach to M2M hyper-parameter determination.
Other than to compare M2M and SCHW orbit classifications for IllustrisTNG data with Xu2019, no external comparisons of models are made.
3 Observational Data
The observational data sets used in this article are the same as those used in Long2025. The data tables here are modified versions of the tables from that paper.
For the ATLAS3D data set (AtlasI), data from four galaxies are utilized - NGC 1248, NGC 3838, NGC 4452, and NGC 4551. These galaxies have various inclinations, different mass-to-light ratios, and sense of bulk rotation (see Table 2). Surface brightness values are calculated using multi-Gaussian expansions (Sect. 4.1). The IFU cells for kinematical data (first and second velocity moments) are the Voronoi cells resulting from signal-to-noise processing of the raw data.
Table 2: Top level data for the ATLAS3D galaxies
Galaxy Inclination M/L ratio Type Bulk rotation SB max radius IFU cells
NGC 1248 $42^{\circ}$ $2.50$ S0 Counter clockwise 3.00 297
NGC 3838 $79^{\circ}$ $4.00$ S0 Clockwise 5.86 383
NGC 4452 $88^{\circ}$ $5.20$ S0 Clockwise 5.17 489
NGC 4551 $63^{\circ}$ $4.89$ E Counter clockwise 3.41 596
The inclinations are taken from AtlasXV, while the M/L ratios were determined by M2M modeling in Long2016. The SB max radius column gives the overall size of the polar grid used for surface brightness data values. The IFU column gives the number of Voronoi cells for the kinematical data.
For the IllustrisTNG galaxies (Nelson2019), five axisymmetric galaxies (A0490, A1090, A1190, A1290, and A1390) are utilized, and are viewed edge on with the mass-to-light ratio taken as $1.0$ for all galaxies. Surface brightness values are again calculated using multi-Gaussian expansions. The IFU cells for kinematical data (the Gauss-Hermite coefficients $h_{1}$ to $h_{4}$) are constructed to represent a real galaxy’s data (see Table 3).
Table 3: Top level data for the simulated galaxies
Galaxy Bulk rotation SB max radius IFU cells
A0490 Counter clockwise 7.00 220
A1090 Counter clockwise 6.00 197
A1190 Clockwise 6.00 110
A1290 Clockwise 6.00 173
A1390 Clockwise 7.00 119
As in Table 2, the SB max radius column gives the overall size of the surface brightness polar grid, and the IFU column gives the number of Voronoi cells. Inclinations for all the simulated galaxies are $90^{\circ}$, and mass-to-light ratios are $1.0$.
For both data sets, data values are symmetrized appropriately for an axisymmetric model/potential. The units used are effective radii for length, $10^{7}$ years for time, and mass in units of the solar mass with luminosity similarly so.
4 Methods
With the exception of M2M, all the methods and techniques in this section are very similar to those in Long2025 and are not described in detail here. The M2M implementation was used recently in Long2021, and is described more fully in Long2016 and other papers by the lead author prior to that.
4.1 Gravitational Potentials
The gravitational potentials and their derivatives used for modeling are derived from MGEs of the surface brightness data (Emsellem1994; Cappellari2002). For the IllustrisTNG simulated data, dark matter potentials are included as spherical Gaussians, and no black hole modeling takes place. For the ATLAS3D data, no dark matter is included for consistency with AtlasXV, and central black holes are modeled as point sources.
4.2 Orbit Initial Conditions
Initial phase space coordinates for stellar orbits are created using the 3I scheme (theoretically based) and the MDJV scheme (observationally based). These initial coordinates are given an overall sense of rotation matching that shown by the observed mean line-of-sight velocity data, with 75% of the orbits having the same sense of rotation and 25% counter-rotating. In general, $8000$ orbits are used for most of the modeling of the observed data sets. For ATLAS3D galaxies NGC 3838 and NGC 4452, $16000$ orbits are used instead so that the elapsed time impact on the SCHW NNLS process, in particular, may be assessed.
4.3 Orbit Classifications
The circularity measure $\lambda_{z}$ (Zhu2018) is used to classify orbits into one of hot ($\left|\lambda_{z}\right|<=0.25$), cold ($\lambda_{z}>=0.80$), warm ($0.25<\lambda_{z}<0.8$) or counter-rotating. The total orbit weights in each of the classes are taken as the overall orbit classification for a galaxy. Comparing orbit classifications produced by M2M and SCHW is a key part of our modeling analysis. Note that orbit classifications are not used as constraints themselves, as in Long2025. However, the sum of the weights is constrained to be equal to one, and so the sum of the orbit classes must also equal one.
4.4 SCHW Models
SCHW has three main phases in its execution.
Phase 1
Using the initial conditions as orbit start coordinates, orbit trajectories are created in the gravitational potential that has been associated with the observed data. The phase space coordinates representing the trajectories are used to construct the individual orbits’ contributions to the model equivalents of the observed data. Phase 1 completes once a predetermined number of orbit integration time steps or orbits have been completed.
Phase 2
The individual orbit contributions are weighted in a non-negative least squares sense (NNLS) such that the model observables reproduce as closely as possible the observed data. Any regularization constraints are included in this phase.
Phase 3
The weighted orbits are analyzed as necessary to learn more about the observed data.
For use with SCHW, the observed data must be modified so that it is surface luminosity (or mass) weighted. In addition, for this investigation, regularization is used (heavy weights penalization) in order to make realistic comparisons with M2M possible (see Sections 4.5 and 5.1).
No use of SCHW dithering (orbit bundling) is made as it causes a significant performance overhead for SCHW.
4.5 M2M Models
M2M differs from SCHW in that phase 1 and phase 2 are combined into a single phase in which the weights exist from the start of modeling and are iteratively adapted, as the orbit trajectories are traversed, such that the model observables (calculated using the weights) eventually reproduce the observed data. This iterative adaption of the weights is akin to the back-propagation mechanism in machine learning neural networks. M2M terminates in a similar fashion to SCHW phase 1. On termination, an extra condition must be met which is that the weights should have converged to constant values (to within some numerical tolerance).
Exponential smoothing is used by default in the construction of M2M model observables. This smoothing mechanism could be turned off but, for the low number of orbits used here, results in much noise within the model, for example in the observable $\chi^{2}$ series. The preferred approach here is to regularize SCHW instead (see Sections 4.5 and 5.1) to achieve an approximate balance in behavior of the two methods.
4.6 Hyper-parameters using Gaussian Processes
Both modeling methods require a number of hyper-parameters to be set, for example, the $\epsilon$ learning rate parameter in M2M, or the regularization parameter in SCHW. In general, these parameters have been set by a combination of experience and experimentation. During this investigation, an experimental routine using Gaussian processes was developed using the scikit-optimize package’s gp_minimize interface. This routine was used to good effect to confirm some of the M2M $\epsilon$ settings, and also to demonstrate that M2M can be used successfully inside a more general framework application.
4.7 Software Utilization
As in Long2025, the software base for constructing initial conditions and for M2M and SCHW modeling is the lead author’s implementation of the Syer1996 M2M stellar dynamical modeling method together with the Schwarz1979 orbit superposition method (a single code set supports both methods). This software was first used in Long2010, and, as indicated, more recently in Long2025. The software is Python 3 based with some use of Cython (behnel2010cython) and C for performance critical code. The MPI message passing interface is used to achieve parallelization of orbit integration, with the same number of cores being used for both M2M and SCHW. The optimization software used for NNLS is taken from the SCIPY package version 1.16 (2020SciPy-NMeth) (earlier versions may have NNLS performance issues). The Gaussian processes application is based on routines within the publicly available scikit-optimize package. Using a $20$ core desktop PC, multi-processor working is employed with up to 10 cores active in a modeling run.
5 Results
Most of the column headings in the tables within this section should be self-explanatory. However, for clarity,
  • $\bullet$
    ‘Active orbits’ refers to the number of orbits that have a non-zero weight while ‘Orbits’ indicates the number of orbits in the initial conditions.
  • $\bullet$
    Under ‘Mean $\chi^{2}$ values’, ‘sb los v1 los v2’ indicates surface brightness and the line-of-sight first and second velocity moments. ‘sb $h_{1}$ $h_{2}$ $h_{3}$ $h_{4}$’ indicates surface brightness and the Gauss-Hermite coefficients $h_{1}$ to $h_{4}$.
  • $\bullet$
    Under ‘Orbit classification’, ‘Hot Warm Cold C-R’ refers to the hot, warm, cold and counter-rotating class values described in Section 4.3.
As an aid to the reader, tables used in comparing orbit classification results employ colored text to indicate differences of $\geq 3\%$ (as in Long2025). What is important in these tables is to be clear on the direction the comparison is being made (horizontally or vertically between sections in the tables), and to use the colors to gain a visual impression of the extent of any disagreement. For example, a table section with few colored entries means few differences implying results are similar; conversely, many colored entries imply that results are dissimilar. Note that the sum of the orbit classes in a classification for a galaxy, as stated in Section 4.3, must be one. This means that an increase or decrease in one class value must result in decreases or increases in one or more other class values.
5.1 Regularizing SCHW
In order to ensure a more balanced comparison between M2M and SCHW, a regularized form of SCHW is used in which heavy weights are penalized. This has a comparable end effect to the exponential smoothing used in M2M, and helps to ensure that the orbit weights of SCHW and M2M in integral space are smoothed to a similar degree. This is illustrated in Figure 1 using models created with the ATLAS3D data for NGC 4452. The other ATLAS3D galaxies and the IllustrisTNG galaxies behave similarly. Smoothness in Figure 1 is calculated as the second derivative of the weights with respect to energy, and to the z-component of angular momentum. As may be seen from the figure, the unpenalized SCHW plots are considerably less smooth than the penalized versions, with penalized SCHW exhibiting a similar degree of smoothness to M2M. In addition, the impact of heavy weights penalization is to bring the SCHW weight range closer to that of M2M, and also to increase the number of active orbits in SCHW. For NGC 4452, the M2M weight range is $(7.09\times 10^{-7},1.71\times 10{-3})$ and the number of active orbits is $15963$. For SCHW, the range is $(2.35\times 10^{-6},5.02\times 10^{-2})$ with $635$ active orbits without penalization, and $(3.03\times 10^{-9},1.01\times 10^{-3})$ and $13106$ active orbits with penalization. These points provide the background to our choice of penalized SCHW for M2M versus SCHW comparisons.
Figure 1: Heavy Weights Penalization illustrated with NGC 4452
SCHW with no penalization SCHW penalized M2M
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Heavy weights penalization illustrated with NGC 4452 (Section 5.1). Smoothness is calculated as the second derivative of the weights with respect to the two integrals ($E$ and $L_{z}$). As may be seen from the figure, the unpenalized SCHW plots show considerably more variation than penalized SCHW. Penalized SCHW, using a parameter value of $1.6\times 10^{3}$, shows a similar degree of smoothness to M2M.
5.2 $\chi^{2}$ and Elapsed Time Comparisons between M2M and SCHW
The focus in this section is on how well observed data are reproduced, and comparing modeling elapsed times. Orbit classifications are considered in the next section. M2M and SCHW models using both 3I and MDJV initial conditions and both data sets are created. Table 4 contains the results for the ATLAS3D galaxy data, and Table 5 for the IllustrisTNG galaxy data. In the tables, all the mean observable $\chi^{2}$ values are $\approx 1$ or less indicating that the observed data have been well-reproduced by both M2M and SCHW. An example is shown in Appendix A. M2M weight convergence values are $93\%$ implying that convergence to constant weights, as expected in Syer1996, is close to being achieved. For ATLAS3D , the modeling elapsed times show that M2M performs faster than SCHW by a factor of $\approx 4$ for 8000 orbit models and by $\approx 5$ for 16000 orbit models. For IllustrisTNG data, M2M outperforms SCHW by a factor of $\approx 3$. Any of these performance improvement factors would benefit large surveys where there are thousands of galaxy data sets to be processed. The NNLS contributions to SCHW elapsed times show that, with increasing model size (numbers of orbits in this case), the performance difference will become more marked depending on the algorithms within the NNLS implementation.
Table 4: ATLAS3D M2M and SCHW 3I and MDJV Comparisons - $\chi 2$ and Elapsed Times
Model Mean $\chi^{2}$ values M2M Weight Elapsed
Galaxy Orbits Duration sb los v1 los v2 Convergence Time
(hmdtu) (s)
3I- M2M
NGC 1248 8000 200 0.31 0.27 0.63 99% 358
NGC 3838 16000 500 1.07 0.18 0.57 96% 1065
NGC 4452 16000 200 0.76 0.42 0.69 96% 865
NGC 4551 8000 200 0.41 0.19 0.76 100% 217
3I- SCHW
NGC 1248 8000 200 0.07 0.11 0.28 1407 (228)
NGC 3838 16000 500 0.19 0.09 0.18 5024 (1379)
NGC 4452 16000 200 0.30 0.13 0.25 4555 (1524)
NGC 4551 8000 200 0.06 0.07 0.23 900 (238)
MDJV - M2M
NGC 1248 8000 200 1.06 0.11 0.30 96% 561
NGC 3838 16000 500 0.91 0.35 0.33 93% 1229
NGC 4452 16000 200 1.08 0.52 1.12 94% 1008
NGC 4551 8000 200 0.52 0.16 1.33 98% 252
MDJV - SCHW
NGC 1248 8000 200 0.05 0.12 0.36 1486 (225)
NGC 3838 16000 500 0.12 0.13 0.21 5422 (1392)
NGC 4452 16000 200 0.56 0.23 0.44 4763 (1422)
NGC 4551 8000 200 0.10 0.07 0.32 1012 (267)
$\chi 2$ and elapsed time M2M and SCHW comparison using ATLAS3D data (Section 5.2). All the mean observable $\chi^{2}$ values are $\approx 1$ or less indicating that the observed data have been well-reproduced by both M2M and SCHW. The M2M weight convergence is $90\%$. The SCHW $\chi^{2}$ values are less than those for M2M due to over-fitting of the observed data (as expected). M2M performs faster than SCHW in elapsed times by a factor of $\approx 4$ for 8000 orbit models and by $\approx 5$ for 16000 orbit models. The numbers in brackets are the NNLS contributions to the elapsed times. With increasing numbers of orbits, the performance difference will become more marked depending on the NNLS implementation used by SCHW.
Table 5: IllustrisTNG M2M and SCHW 3I and MDJV Comparisons - $\chi 2$ and Elapsed Times
Model Mean $\chi^{2}$ values M2M Weight Elapsed
Galaxy Orbits Duration sb $h_{1}$ $h_{2}$ $h_{3}$ $h_{4}$ Convergence Time
(hmdtu) (s)
3I - M2M
A0490 8000 200 0.94 0.72 1.10 0.66 1.23 96% 1203
A1090 8000 100 0.87 0.93 1.04 0.63 0.61 93% 632
A1190 8000 100 0.75 0.79 0.92 0.84 0.56 97% 738
A1290 8000 100 0.71 1.01 0.94 0.77 0.73 93% 550
A1390 8000 100 0.90 0.85 1.24 0.61 0.56 96% 628
3I - SCHW
A0490 8000 200 1.30 0.39 0.72 0.43 0.89 3746 (196)
A1090 8000 100 0.54 0.49 0.59 0.25 0.31 1948 (200)
A1190 8000 100 0.53 0.35 0.73 0.20 0.39 2240 (197)
A1290 8000 100 0.33 0.37 0.41 0.23 0.38 1816 (200)
A1390 8000 100 0.69 0.42 0.88 0.28 0.35 2037 (200)
MDJV - M2M
A0490 8000 100 1.04 0.75 0.79 0.54 1.07 96% 753
A1090 8000 100 0.78 0.88 1.01 0.52 0.43 95% 776
A1190 8000 100 0.61 0.60 0.58 0.43 0.53 96% 858
A1290 8000 100 1.21 0.70 0.68 0.64 0.54 97% 665
A1390 8000 100 0.85 0.81 1.06 0.46 0.44 96% 776
MDJV - SCHW
A0490 8000 100 0.43 0.20 0.21 0.21 0.42 1939 (235)
A1090 8000 100 0.54 0.31 0.42 0.24 0.20 2180 (257)
A1190 8000 100 0.26 0.22 0.31 0.15 0.21 2336 (267)
A1290 8000 100 0.37 0.29 0.24 0.15 0.21 1865 (258)
A1390 8000 100 0.48 0.33 0.44 0.20 0.25 2193 (258)
$\chi 2$ and elapsed time M2M and SCHW comparison using IllustrisTNG data (Section 5.2). All the observable $\chi^{2}$ values are $\approx 1$ or less indicating that the observed data have been well-reproduced by both M2M and SCHW with, for M2M, the weight convergence being $90\%$. M2M outperforms SCHW by a factor of $\approx 3$ in elapsed times. As in Table 4, the numbers in brackets are the NNLS contributions to the elapsed times.
5.3 Orbit Classification Comparisons between M2M and SCHW
The comparisons between orbit classification results from modeling with M2M and SCHW are shown in Table 6. M2M and SCHW are producing similar orbit classifications from the same observed data and initial conditions, with ATLAS3D and MDJV showing the most variation in classification and IllustrisTNG and MDJV showing the least.
Table 6: M2M versus SCHW Comparisons for 3I and MDJV - Orbit Classifications
3I Initial Conditions MDJV Initial Conditions
Galaxy Active Orbit Classification Active Orbit Classification
Orbits Hot Warm Cold C-R Orbits Hot Warm Cold C-R
Atlas3D – M2M
NGC 1248 7977 0.18 0.41 0.23 0.17 8000 0.24 0.43 0.21 0.12
NGC 3838 14098 0.16 0.37 0.36 0.11 13742 0.27 0.51 0.16 0.06
NGC 4452 15896 0.17 0.31 0.34 0.17 15389 0.25 0.49 0.12 0.14
NGC 4551 7889 0.26 0.39 0.15 0.19 7882 0.39 0.46 0.03 0.13
Atlas3D – SCHW
NGC 1248 $78$ 0.19 0.42 0.25 0.13 7804 0.25 0.45 0.17 0.13
NGC 3838 10484 0.22 0.39 0.31 0.07 12440 0.23 0.56 0.15 0.07
NGC 4452 13106 0.19 0.31 0.31 0.18 8678 0.19 0.51 0.13 0.17
NGC 4551 7771 0.30 0.38 0.13 0.18 7225 0.33 0.48 0.03 0.16
IllustrisTNG – M2M
A0490 8000 0.36 0.28 0.13 0.23 8000 0.36 0.34 0.09 0.21
A1090 8000 0.30 0.43 0.15 0.12 8000 0.34 0.46 0.09 0.11
A1190 8000 0.39 0.40 0.08 0.13 8000 0.42 0.41 0.04 0.13
A1290 8000 0.31 0.34 0.20 0.14 7977 0.34 0.37 0.15 0.14
A1390 8000 0.33 0.42 0.12 0.13 8000 0.34 0.46 0.06 0.13
IllustrisTNG - SCHW
A0490 4529 0.38 0.31 0.14 0.18 6084 0.37 0.34 0.09 0.20
A1090 5315 0.30 0.41 0.18 0.11 6906 0.35 0.46 0.10 0.10
A1190 5581 0.40 0.37 0.10 0.13 7379 0.43 0.40 0.04 0.13
A1290 5135 0.30 0.36 0.21 0.13 6984 0.34 0.38 0.15 0.13
A1390 5739 0.33 0.40 0.13 0.13 7219 0.35 0.46 0.06 0.12
Orbit classification comparisons for M2M and SCHW with both 3I and MDJV initial conditions (see Section 5.3). Comparisons in the table are made vertically, first between M2M and SCHW using ATLAS3D data, and then using IllustrisTNG data. Classification differences of $\geq 3\%$ between M2M and SCHW are indicated in blue. M2M and SCHW are producing similar orbit classifications from the same observed data, with ATLAS3D and MDJV showing the most variation (top right ‘partition’) and IllustrisTNG and MDJV showing the least (bottom right ‘partition’). The largest individual variation is $6\%$ and occurs in the hot classifications for ATLAS3D data.
Orbit classifications are compiled from individual orbits and, given that here the same sets of orbits (initial conditions) are used for M2M and SCHW modeling, whether or not M2M and SCHW are using the same orbits in the same classifications is investigated. Table 7 shows how many orbits are common to both modeling methods, and how many are specific to one of M2M or SCHW. Comparatively few are specific to SCHW only, with substantially more being specific to M2M. This is readily explainable: orbits which are zero weighted by SCHW appear in the M2M only figures.
Table 7: M2M versus SCHW Orbit Comparison
3I Initial Conditions MDJV Initial Conditions
Galaxy Active Orbits Common M2M SCHW Active Orbits Common M2M SCHW
M2M SCHW Orbits only only M2M SCHW Orbits only only
ATLAS3D
NGC 1248 7992 7886 7843 149 3 7985 7812 7759 226 9
NGC 3838 15947 10528 10087 4992 290 15763 12475 12241 2645 113
NGC 4452 15963 13106 12910 3053 20 15992 8678 8587 7251 7
NGC 4551 7975 7780 7726 232 22 7955 7225 7158 797 7
IllustrisTNG
A0490 7993 4563 4519 3474 0 8000 6087 6032 1968 0
A1090 8000 5315 5259 2741 0 8000 6906 6855 1145 0
A1190 8000 5581 5523 2416 0 8000 7390 7339 661 0
A1290 8000 5135 5074 2926 0 8000 6992 6926 1045 11
A1390 8000 5739 5684 2316 0 8000 7219 7165 823 0
M2M versus SCHW orbit comparisons (Section 5.3) using both sets of initial conditions and both data sets. The comparison for every galaxy’s data shows the number of orbits which are common to both M2M and SCHW models, and those that are particular to just one model. Zero-weighted orbits in SCHW models, if they are used by M2M, are accounted for in the M2M only subset. This explains the comparatively small values in the SCHW only column, and the much larger values in the M2M only column. The common orbits are examined in more detail in Figure 2.
The next step is to look in detail at the common orbits, to investigate whether individual orbits are weighted similarly by M2M and SCHW. Figure 2 contains various plots from considering just the ATLAS3D data set modeled with 3I initial conditions. The left hand scatter plots show the paired orbit weights arising from M2M and SCHW modeling, with the blue dashed line indicating equality of the orbit weights. The key point to note is that the weights are not tightly positioned close to the equality line: even though the orbits used are the same, M2M and SCHW are in general not allocating the same weights to the orbits. The central Bland-Altman plots of the average log weight for an orbit against the log difference between its M2M weight and its SCHW weight show that the mean log difference (red solid line) is approximately zero (absolute values $10^{-2}$) indicating that there is no significant bias towards either method. In addition, weight differences are larger at lower average weights than with higher average weights. The red dashed lines show the $2.5\%$ and $97.5\%$ percentiles with $95\%$ of the difference ratios lying between the lines. Expressing the values of the percentiles as percentages, they are equivalent to saying that the weight differences range from a few tens of percent lower to a few hundred percent higher. The right hand plots show the distributions of the weight ratios. The red lines have the same meaning as in the Bland-Altman plots.
Figure 2: M2M and SCHW Common Orbits Weight Comparison for ATLAS3D Data
M2M vs SCHW Weights Bland-Altman Log Weights Ratio
NGC 1248 Refer to caption Refer to caption Refer to caption
NGC 3838 Refer to caption Refer to caption Refer to caption
NGC 4452 Refer to caption Refer to caption Refer to caption
NGC 4551 Refer to caption Refer to caption Refer to caption
M2M and SCHW common orbits weight comparisons for ATLAS3D data (Section 5.3). 3I initial conditions are used for the models. The left hand column shows scatter plots of the weights of the orbits, with the blue dashed line indicating equality of the orbit weights. The key point to note is that the weights are not tightly positioned close to the equality line: even though the orbits used are the same, M2M and SCHW are in general not allocating the same weights to the orbits. The Bland-Altman plots (BA1983) of the average log weight for an orbit against the log difference between its M2M weight and its SCHW weight show that the mean log difference (red solid line) is approximately zero indicating that there is no significant bias towards either method. In addition, weight difference is larger at lower average weights than with larger average weights. The red dashed lines show the $2.5\%$ and $97.5\%$ percentiles with $95\%$ of the difference ratios lying between the lines. The right hand plots show the distributions of the weight ratios. The red lines have the same meaning as in the Bland-Altman plots.
Finally, it is useful to note that if in a model orbit weights are sorted into descending order, then just how many orbits contribute to the heaviest $25\%$, $50\%$, $75\%$ and $100\%$ by weight can be established. Using 3I initial conditions, the approximate profile for ATLAS3D is $8\%$, $22\%$, $44\%$, $100\%$] of the orbits, and for IllustrisTNG $4\%$, $14\%$, $35\%$, $100\%$]. What is then clear is the first $25\%$ of the total weight involves a small number of heavily weighted orbits, while over $50\%$ orbits (lightly weighted) make up the final $25\%$. Both M2M and SCHW are affected in the same way.
5.4 Comparisons between 3I and MDJV
The orbit classification results achieved with 3I and MDJV initial conditions are compared for M2M and SCHW using both 3I and MDJV initial conditions. The comparisons are shown in Table 8. The numerical values in Table 8 are the same as Table 6 but this time the comparisons are made horizontally (3I versus MDJV) rather than vertically (M2M versus SCHW). From the coloring used, it should be apparent from Table 8 that 3I and MDJV initial conditions do not produce the same orbit classifications. The only class where there is some agreement is with the counter-rotating orbits.
Table 8: 3I versus MDJV Comparisons for M2M and SCHW - Orbit Classifications
3I Initial Conditions MDJV Initial Conditions
Galaxy Active Orbit Classification Active Orbit Classification
Orbits Hot Warm Cold C-R Orbits Hot Warm Cold C-R
Atlas3D – M2M
NGC 1248 7977 0.18 0.41 0.23 0.17 8000 0.24 0.43 0.21 0.12
NGC 3838 14098 0.16 0.37 0.36 0.11 13742 0.27 0.51 0.16 0.06
NGC 4452 15896 0.17 0.31 0.34 0.17 15389 0.25 0.49 0.12 0.14
NGC 4551 7889 0.26 0.39 0.15 0.19 7882 0.39 0.46 0.03 0.13
Atlas3D – SCHW
NGC 1248 $78$ 0.19 0.42 0.25 0.13 7804 0.25 0.45 0.17 0.13
NGC 3838 10484 0.22 0.39 0.31 0.07 12440 0.23 0.56 0.15 0.07
NGC 4452 13106 0.19 0.31 0.31 0.18 8678 0.19 0.51 0.13 0.17
NGC 4551 7771 0.30 0.38 0.13 0.18 7225 0.33 0.48 0.03 0.16
IllustrisTNG – M2M
A0490 8000 0.36 0.28 0.13 0.23 8000 0.36 0.34 0.09 0.21
A1090 8000 0.30 0.43 0.15 0.12 8000 0.34 0.46 0.09 0.11
A1190 8000 0.39 0.40 0.08 0.13 8000 0.42 0.41 0.04 0.13
A1290 8000 0.31 0.34 0.20 0.14 7977 0.34 0.37 0.15 0.14
A1390 8000 0.33 0.42 0.12 0.13 8000 0.34 0.46 0.06 0.13
IllustrisTNG - SCHW
A0490 4529 0.38 0.31 0.14 0.18 6084 0.37 0.34 0.09 0.20
A1090 5315 0.30 0.41 0.18 0.11 6906 0.35 0.46 0.10 0.10
A1190 5581 0.40 0.37 0.10 0.13 7379 0.43 0.40 0.04 0.13
A1290 5135 0.30 0.36 0.21 0.13 6984 0.34 0.38 0.15 0.13
A1390 5739 0.33 0.40 0.13 0.13 7219 0.35 0.46 0.06 0.12
The orbit classification comparisons for 3I and MDJV initial conditions (Section 5.4). Comparisons in the table are made horizontally with classification differences of $\geq 3\%$ between 3I and MDJV being indicated in red. 3I and MDJV do not produce the same orbit classifications, though it is interesting that there is less disagreement on the counter-rotating orbit classification.
5.5 Comparison with Xu2019
The orbit classification comparisons for M2M and SCHW using IllustrisTNG data in Section 5.3 are compared with Xu2019, and the results recorded in Table 9. Apart from the A0490 M2M model, all the other models have differences with Xu2019 in the hot, warm and cold classifications and this is considered further in the Discussion (Section 6). Apart from the counter-rotating classification for the A0490 SCHW model, all other counter-rotating classifications agree, to within $3\%$, with Xu2019. The differences in the MDJV cold column reflect the fact that MDJV produces fewer cold orbits than 3I, and does not meet the orbital needs of the observed data.
Table 9: IllustrisTNG M2M and SCHW Comparisons with Xu2019 - Orbit Classifications
3I Initial Conditions Xu2019 MDJV Initial Conditions
Galaxy Orbit Classification Orbit Classification Orbit Classification
Hot Warm Cold C-R Hot Warm Cold C-R Hot Warm Cold C-R
IllustrisTNG – M2M
A0490 0.36 0.28 0.13 0.23 0.35 0.28 0.14 0.23 0.36 0.34 0.09 0.21
A1090 0.30 0.43 0.15 0.12 0.28 0.46 0.15 0.10 0.34 0.46 0.09 0.11
A1190 0.39 0.40 0.08 0.13 0.33 0.43 0.11 0.12 0.42 0.41 0.04 0.13
A1290 0.31 0.34 0.20 0.14 0.32 0.37 0.18 0.13 0.34 0.37 0.15 0.14
A1390 0.33 0.42 0.12 0.13 0.29 0.45 0.15 0.12 0.34 0.46 0.06 0.13
IllustrisTNG - SCHW
A0490 0.38 0.31 0.14 0.18 0.35 0.28 0.14 0.23 0.37 0.34 0.09 0.20
A1090 0.30 0.41 0.18 0.11 0.28 0.46 0.15 0.10 0.35 0.46 0.10 0.10
A1190 0.40 0.37 0.10 0.13 0.33 0.43 0.11 0.12 0.43 0.40 0.04 0.13
A1290 0.30 0.36 0.21 0.13 0.32 0.37 0.18 0.13 0.34 0.38 0.15 0.13
A1390 0.33 0.40 0.13 0.13 0.29 0.45 0.15 0.12 0.35 0.46 0.06 0.12
Orbit classification comparisons for M2M and SCHW with Xu2019 (see Section 5.5). Comparisons in the table are made horizontally between models and Xu2019. Classification differences of $\geq 3\%$ with Xu2019 are indicated in red and blue. Apart from the A0490 M2M model, all the other models have some differences with Xu2019. The values in the counter-rotating columns are notable for their lack of differences. The MDJV cold column reflects differences between the initial conditions where MDJV has fewer cold orbits than 3I.
5.6 Impacts of Changing Data Layout
In this section, the impact of interpolating the IFU kinematical data from their Voronoi cells onto a polar grid is considered. The polar grid chosen has 8 divisions logarithmically in radius and 32 in angle (making up $2\pi$ radians). The Voronoi observed data are interpolated onto this $(8,32)$ polar grid using ‘nearest point’ interpolation (see Figure 3). The modeling runs are limited to M2M with 3I initial conditions and ATLAS3D galaxy data only: no use is made of SCHW, MDJV or IllustrisTNG data for this part of the investigation.
Figure 3: Data Positions for NGC 4452
Refer to caption Refer to caption
IFU and polar data positions for NGC 4452 kinematical data (Section 5.6). The left hand plot shows the IFU data positions in yellow with the black dashed line showing the convex hull enclosing the data points. The blue circle has the minimum radius necessary to surround all data points. This radius gives the extent of the polar grid in the right hand plot. In this plot, the yellow data points have data values interpolated from the IFU data values. The red points are the points in the polar grid that are outside of the convex and have no data values.
No issues with observable $\chi^{2}$ values were noted and the M2M weights converged as required (Table 10).
Table 10: Voronoi and Polar Kinematical Data - $\chi 2$ Values
Mean $\chi^{2}$ values M2M Weight
Galaxy Orbits sb los v1 los v2 Convergence
Polar kinematical data
NGC 1248 8000 0.38 0.23 0.25 99%
NGC 3838 16000 1.01 0.51 0.50 92%
NGC 4452 16000 1.16 0.31 1.07 96%
NGC 4551 8000 0.87 0.24 1.17 98%
Voronoi kinematical data
NGC 1248 8000 0.31 0.27 0.63 99%
NGC 3838 16000 1.07 0.18 0.57 96%
NGC 4452 16000 0.76 0.42 0.69 96%
NGC 4551 8000 0.41 0.19 0.76 100%
The observable mean $\chi^{2}$ values from the M2M ATLAS3D polar versus Voronoi cells comparison (Section 5.6). All the $\chi^{2}$ values are $\approx 1$ or less, and the weight convergence values are $90\%$ indicting that M2M has reproduced the observed data as expected.
Orbit classification values are in Table 11. Nine differences $\geq 3\%$ are recorded, with the maximum difference being $9\%$. Overall, it appears that the geometric arrangement of the observed data does influence the orbit classifications obtained. Given the results obtained in the investigation as a whole, it is considered unlikely that the geometric effect is M2M specific and does not affect SCHW as well.
Table 11: Voronoi and Polar Kinematics - Orbit Classification
Active Orbit Classification
Galaxy Orbits Orbits Hot Warm Cold C-R
Polar kinematics
NGC 1248 8000 7782 0.21 0.41 0.23 0.15
NGC 3838 16000 12749 0.23 0.40 0.27 0.11
NGC 4452 16000 16000 0.18 0.28 0.28 0.25
NGC 4551 8000 7824 0.29 0.41 0.12 0.18
Voronoi kinematics
NGC 1248 8000 7977 0.18 0.41 0.23 0.17
NGC 3838 16000 14098 0.16 0.37 0.36 0.11
NGC 4452 16000 15896 0.17 0.31 0.34 0.17
NGC 4551 8000 7889 0.26 0.39 0.15 0.19
The orbit classification values from the M2M ATLAS3D polar versus Voronoi cells comparison (Section 5.6). Table values in red indicate a classification difference of $\geq 3\%$. In total, there are nine red values suggesting that orbit classifications are dependent on the geometric arrangement of the observed data. The largest difference in classification is $9\%$.
The impact of changing the surface brightness layout is now briefly considered. Taking one galaxy’s ATLAS3D data (NGC 4551), 3I initial conditions, and modeling with M2M, the following polar grid layouts are employed: $(16,8)$ and $(8,16)$ with uniform radial intervals, and $(16,16)$ with uniform and logarithmic radial intervals. This gives four different surface brightness layouts. The kinematical data are unmodified from the IFU data used for the NGC 4551 3I modeling run in Table 4. The results obtained for the four modeling runs are
  1. 1.
    the mean $\chi^{2}$ values for surface brightness range from $0.14$ to $0.41$,
  2. 2.
    the $\chi^{2}$ values for the first and second velocity moments are the same for all runs,
  3. 3.
    M2M weight convergence is $99\%$, and
  4. 4.
    the orbit classifications are the same for all runs to within $1\%$.
In short, varying the surface brightness polar layout in isolation has had little effect on the modeling results achieved. This clearly differs from the kinematical data results described above.
6 Discussion
The Results (Section 5) contain many tables and much detail so the main findings are summarized below.
  1. 1.
    M2M and SCHW consistently achieve mean observable $\chi^{2}\approx 1$ or less for both data sets and initial conditions. Related, M2M weight convergence is $90\%$. In other words, both methods are able to reproduce observed data to an acknowledged, acceptable level (see Tables 4 and 5).
  2. 2.
    In terms of computer elapsed time, using the same computer configuration, M2M consistently outperforms SCHW. M2M elapsed times are between three and five times faster than SCHW (see Tables 4 and 5).
  3. 3.
    For both data sets, M2M and SCHW produce orbit classifications which are approximately the same for the same initial conditions (Table 6), but are not the same for different initial conditions (Table 8). The classifications are dependent on the initial conditions.
  4. 4.
    Both M2M and SCHW using IllustrisTNG data as in Xu2019 do not produce orbit classifications which match those in Xu2019 (see Table 9). It therefore appears that the gravitational potential, observed data sets and initial conditions used by the methods are not able to produce good ‘orbit level’ representations of the associated galaxies.
  5. 5.
    M2M and SCHW, using the same gravitational potentials and initial conditions, do not use exactly the same orbits in reproducing the same observed data. For orbits which are common, these orbits do not have the same weights (see Figure 2).
  6. 6.
    Changing the data layout of surface brightness from one polar system to another does not appear to affect model results. However, changing the kinematical data layout from IFU Voroni cells to a polar system does cause the orbit classification results to change (see Table 11).
Nothing has come out of the investigation which indicates that the performance benefit that comes from using M2M should not be taken advantage of. The benefits to galaxy survey modeling are obvious. In addition, M2M is not affected by the performance limitations of using NNLS. Also, if the performance cost of NNLS were to be reduced to close to zero, M2M would still be faster. All that is required is for researchers to continue to be comfortable with modeling schemes using low orbit numbers to reproduce galaxy data, and with creating models that do not necessarily truly represent the parent galaxy and may actually be non-physical (as another well-known scheme may do). Do not forget that, regardless of regularization, orbit selection within the methods is purely numerical and not astrophysical.
All comparisons between M2M and SCHW will, at some level, be implementation specific, but there seems to be little to be gained by seeking to address known modeling shortfalls in an SCHW-like context. M2M is inherently more flexible and scalable than SCHW. Orbit weights are available within the model from the outset so more complex observables are easier to model or estimate. Data uncertainties are an integral part of the M2M weight determination process, and are accommodated far more so than SCHW does in reaching the final solution. Once designed, achieving a basic M2M implementation in approximately six to eight weeks can be accomplished by an experienced programmer using a high level language such as Python, and including use of parallelization. Creating a GPU based implementation for M2M can be achieved using tooling such as PyTorch (paszke2019). However to gain the full performance benefits, this will require careful assessment and design of the GPU parallelizable features of M2M. Given that M2M employs a number of techniques that are used in machine learning, perhaps they might provide an alternative way forward and this will be investigated in the future by the authors
7 Conclusions
The objective set out in the Introduction to assess and compare the operational capabilities of M2M and SCHW has been met. Both methods are able to replicate observed data to an acceptable level but M2M outperforms SCHW being several times faster. This speed improvement makes M2M more desirable in a galaxy survey context than SCHW. Both methods produce similar orbit classifications for a given set of initial conditions, but the classifications do appear to be particular to the initial conditions used. In addition, the weights of the common orbits under-pinning the classifications differ and are method specific. Looking ahead, M2M is the more attractive option currently for stellar dynamical modeling in the 21st century, but could eventually be overtaken by machine learning based methods. However, in the mean time, this paper acts as a flag to the stellar dynamical modeling community that M2M should be seriously considered as a way of alleviating some of the shortcomings of SCHW.
Acknowledgements.
The author thanks Dandan Xu, Ling Zhu, Shude Mao and Yunpeng Jin for various fruitful discussions as the research recorded here progressed.
Appendix A Example Observable Reproduction
This appendix contains a visual example in Figure 4 of the reproduction of observables by M2M and SCHW using NGC 4452 with 3I initial conditions (see Section 5.2).
Figure 4: M2M and SCHW Reproduction of Observables - NGC 4452
Observed Data M2M SCHW
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Reproduction of observables by M2M and SCHW using NGC 4452 with 3I initial conditions (Appendix A). The observables are surface brightness in the top row followed by the first and second velocity moments. Units are as in Section 3. The first column displays the observed data used in modeling, with the second and third columns showing reproduction of the observed data by the modeling methods. In row order, the mean observable $\chi^{2}$ values are ($0.76$, $0.42$, $0.69$) for M2M and ($0.30$, $0.13$, $0.25$) for SCHW (see Table 4).